40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 23 16:41:06 2020, Koji, Update, IOO, Excess laser freq noise investigation plot.pdf
    Reply  Mon Oct 26 13:35:58 2020, Koji, Update, IOO, Excess laser freq noise investigation 6x
       Reply  Mon Oct 26 17:26:26 2020, gautam, Update, IOO, Excess laser freq noise investigation 
Message ID: 15643     Entry time: Mon Oct 26 13:35:58 2020     In reply to: 15641     Reply to this: 15644
Author: Koji 
Type: Update 
Category: IOO 
Subject: Excess laser freq noise investigation 

In fact, the problem was the grounding issue (presumably on the IOO racks).
A temporary differential receiver at the TTFSS side was built using an SR560 and a few ponoma cables. This removed the structures ~850Hz.


The MC Servo Output was disconnected from the TTFSS box and monitored with SR785. The 850Hz structure was kept visible no matter what cables, including all the acromag DB cables, were removed. This made me suspicious about the measurement setup. The SR785 was connected to an AC power strip under the SP table and this was too far from the IOO rack.

The SR785 was connected to the AC power strip on 1X2, and now the difference becomes clear. No matter if the acromag cables are connected or not, the connection (particularly ground connection) between the MC servo module and the TTFSS box causes the MC servo output contaminated. (Comparison between Blue and Orange of Attachment #1). During the measurement, the EPICS switch for the fast path was disengaged (=no signal) and the VCO gain (...so called. It's just the MC Servo Gain) was set to be 0dB.

To test if the differential receiving of the MC Servo Output at the PSL helps to reduce this noise, I've built a simple (hacky) differential receiver using an SR560. (Attachment #2)
This kept the noise level same as the disconnected case (Comparison between Green and Orange of Attachment #1, I don't think the difference between them is not significant), while the IMC is locked as before.
Note that we can see that the 36kHz line was significantly reduced. Did we remove this annoying noise?

After talking with Gautam, we decided to leave this configuration while the SE-Diff cable was replaced with a more robust one. (See Attachment #3)


The PSL laser frequency performance was evakluated in the following two ways as we did last week:
1) Use the beat frequency of the free running PSL and the Y-end laser (Attachment #4). The PSL shutter was closed and thus the IMC was not locked.
2) Use the IMC MCF while the IMC was locked. (Attachment #5)

For both cases, the improvement was confirmed.


I also tried to check the reported issue by Gautam on this elog. He used 1Hz BW, but I cheated with 16Hz BW and 10x12.8kHz span PSDs. (Attachment #6)

For the measurement, IN1 GAIN of the IMC Servo was set to be 0dB and the OUT2 was switched to monitor the IN1 noise, while IN1 was terminated by a 50Ohm.

As I mentioned above, the AC power of SR785 was taken from a 1X2 power strip. Is this the reason for the power line forest look less severe compared to the previous case???
Anyway, I tried to use the same differential receiving technique (but with gain of x100) to see if this helps. The differential receiver helped to reduce the structure above 50kHz. The floor noise level was observed to be higher. I didn't pursue this any further, but the forest of the power line looked like a part of the measurement noise. This is indicative that the grounding condition on 1X2 is really not great and we need to review the configuration of the acromag grounding.

Attachment 1: MC_Servo_Output.pdf  175 kB  Uploaded Mon Oct 26 19:15:20 2020  | Hide | Hide all
MC_Servo_Output.pdf
Attachment 2: 20201026135735_IMG_0175.jpg  156 kB  Uploaded Mon Oct 26 20:17:36 2020  | Hide | Hide all
20201026135735_IMG_0175.jpg
Attachment 3: 20201026153435_IMG_0176.jpg  93 kB  Uploaded Mon Oct 26 20:21:29 2020  | Hide | Hide all
20201026153435_IMG_0176.jpg
Attachment 4: Screen_Shot_2020-10-26_at_1.15.54_PM.png  90 kB  Uploaded Mon Oct 26 20:22:09 2020  | Hide | Hide all
Screen_Shot_2020-10-26_at_1.15.54_PM.png
Attachment 5: Screen_Shot_2020-10-26_at_1.35.19_PM.png  104 kB  Uploaded Mon Oct 26 20:22:09 2020  | Hide | Hide all
Screen_Shot_2020-10-26_at_1.35.19_PM.png
Attachment 6: MC_Servo_Error_Mon.pdf  193 kB  Uploaded Mon Oct 26 20:32:02 2020  | Hide | Hide all
MC_Servo_Error_Mon.pdf
ELOG V3.1.3-